OpenStudy (anonymous):

Can anyone please help and explain? Find the distance between the points (-3,2) and (-7,4)

OpenStudy (anonymous):

To find the distance between the points : \[A(x_A,y_A)~~\text{ and } B(x_A,y_A)\] We have : \[AB=\sqrt{(x_A-x_B)^2+(y_B-y_A)^2}\] Can you use this to answer your question ?

OpenStudy (anonymous):

Thank you very much @Noura11! Is the answer \[2\sqrt{10}\] correct?

OpenStudy (anonymous):

No ! recalculate it ! ;)

OpenStudy (skullpatrol):

@Noura11 Point B has the same subscripts as point A.

OpenStudy (anonymous):

@skullpatrol It is a mistype only ;)

OpenStudy (skullpatrol):

and the order in the AB formula is not right

OpenStudy (anonymous):

The order is not important because : \[(a-b)^2=(b-a)^2\quad\forall a,b\in{\mathbb R}\]

OpenStudy (skullpatrol):

It is when you are trying to LEARN it for the first time.

OpenStudy (anonymous):

Sorry I am confused :(

OpenStudy (skullpatrol):

See^

OpenStudy (anonymous):

@RH I will explain it to you : \[\text{Let } A(-3,2) \text{ and } B(-7,4)\] Then : \[AB=\sqrt{(-7-(-3))^2+(4-2)^2}=\sqrt{(-4)^2+2^2}=\sqrt{16+4}=\sqrt{20}=\sqrt{4\times5}=2\sqrt5\]

OpenStudy (anonymous):

\[AB=\sqrt{20}=\sqrt{4\times5}=2\sqrt5\]

OpenStudy (anonymous):

@Noura11 Thank you so much!!!!!